10. A random sample of 100 pumpkins is obtained and the mean circumference is found to be 40.5cm. Assuming that the population standard deviation is known to be 1.6cm, use a 0.05 significance level to test the claim that the mean circumference of all pumpkins is equal to 39.9cm



1
Expert's answer
2021-12-27T04:13:19-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=39.9"

"H_0:\mu\not=39.9"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{40.5-39.9}{1.6/\sqrt{100}}=3.75"

Since it is observed that "|z| = 3.75 > 1.96=z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z>3.75)=0.000177," and since "p=0.000177<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is evidence to claim that the population mean "\mu"

is different than "39.9," at the "\alpha = 0.05" significance level.